Sandra Baker held a $3,500, 70-day note dated March 14, bearing interest at 10%. On April 14, she took the note to Loyal Capital Bank which discounted it at 15%. Use a 365-day year to find the missing information on the loan.
Answer #1

Days of maturity remaining on April 14 for note = 70 - 31 = 39 days

Principal, P = $ 3,500; Discount rate, R = 15%; Time, T = 39 / 365 year

Missing information:

Interest amount on the loan = P x R x T = $3,500 x 15% x 39/365 = $ 56.10

Amount paid by Loyal Capital Bank to Sandra = P - interest amount = 3,500 - 56.10 = $ 3,443.90